1 4.6.3. (Harder!) Let 0 < a < 1. Prove that for any n EN, (1 – a)” < 1+n·a
Answer #1

Sol. 6t Osaal then for DEN 2 (-1) (-n-) + (1+xJ, -nx + 2 bolal -a ot and 2 = o < a <I 2 then n (0+1) 36 + (1-a) = 1+ na q